# The lower $p$-series of analytic pro-$p$ groups and Hausdorff dimension

### v2

@@ 2. Stratifications of the lower $p$-series
$[P_{i-2cd}(G),P_{i-2cd}(G)]\subseteq P_{2i-4cd}(G)\subseteq P_{i+2cd}(G).$
In conjunction with [6, Cor. 3.5], the description of the Lie bracket in terms of the group multiplication, as described in [5, Sec. 4] and recorded in [15, Equ. (2.1)] yields the corresponding result on the Lie lattice side,
$[L_{i^{*}-2cd},L_{i^{*}-2cd}]_{\mathrm{Lie}}\subseteq L_{i^{*}+2cd},$
where we write $[\cdot,\cdot]_{\mathrm{Lie}}$ for the Lie bracket to distinguish it from group commutators. Similar to the simpler situation discussed above, the finite abelian group section $P_{i-2cd}(G)/P_{i+2cd}(G)$ of $G$ corresponds to a finite abelian Lie ring section $L_{i^{*}-2cd}/L_{i^{*}+2cd}$ of the $\mathbb{Z}_{p}$-Lie lattice $L$: the multiplicative cosets are the same as the additive cosets, multiplication and addition are the same and the Lie bracket is trivial; compare [8, Thm. 4.5].
Indeed, we already observed that the Lie bracket on $L_{i^{*}-2cd}$ is trivial modulo $L_{i^{*}+2cd}$, and the Lie ring $L_{i^{*}-2cd}/L_{i^{*}+2cd}$ is abelian. In conjunction with [15, Lem. B.2], the description of group multiplication via the Hausdorff series $\Phi$, namely
@@
where $\mu_{U}$ denotes the normalised Haar measure on the group $U$. We recall that the normalised Haar measures $\mu_{U}$ and $\mu_{L}$ on the compact space $U=L$ are the same; see [15, Prop. A.2] and [4, Lem. 1.9.4]. This gives
@@
The uniqueness of the parameters $\xi_{1},\ldots,\xi_{d}$ follows by looking at the elementary divisors of increasing sections, similar to the proof of Lemma 2.1. Indeed, suppose that $j$, $U$, $\xi_{1},\ldots,\xi_{d}$, $c$ and $\check{U}_{i}$ satisfy the assertions of Theorem 1.2. Let $L$ be the $\mathbb{Z}_{p}$-Lie lattice associated to $U$. By [5, Prop. 3.9], there is $j^{\prime}\in\mathbb{N}$ with $j^{\prime}\geq j$ such that, for all $i\in\mathbb{N}$ with $i\geq j^{\prime}$, the subgroups $P_{i}(G),\check{U}_{i}\trianglelefteq_{\mathrm{o}}G$ are uniformly powerful and thus give rise to full $\mathbb{Z}_{p}$-sublattices $A_{i},\check{M}_{i}\leq_{\mathrm{o}}\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$, by the Lie correspondence.
@@ 3. Hausdorff spectra with respect to the lower $p$-series
Furthermore, every subgroup $H\leq_{\mathrm{c}}L$ has strong Hausdorff dimension with respect to $\mathcal{S}$.
@@
We write $\mathcal{S}\colon L=L_{0}\geq L_{1}\geq\ldots$ and choose a frame $(x_{1},\ldots,x_{d})$ for a stratification of $\mathcal{S}$. We also recall that $0<\xi_{1}\leq\xi_{2}\leq\ldots\leq\xi_{d}$. By [17, Lem. 2.2], any two equivalent filtration series define the same Hausdorff dimension on closed subgroups of $L$. Hence we may suppose that
@@
Let $H\leq_{\mathrm{c}}L$ be a $\mathbb{Z}_{p}$-sublattice of dimension $\dim_{\mathbb{Z}_{p}}(H)=r$, say. Since $\mathbb{Z}_{p}$ is a principal ideal domain, $H$ has a $\mathbb{Z}_{p}$-basis $(y_{1},\ldots,y_{r})$ such that the parameters
@@
It suffices to prove that the upper limit on the right-hand side equals 00; in particular, this also shows that it is a proper limit. Since $H_{2}=\mathbb{Z}_{p}y_{r}$, we find that for each $i\in\mathbb{N}$ there exists $n_{i}\in\mathbb{N}_{0}$ such that
@@
As in the proof of Theorem 1.2 we find $j\in\mathbb{N}$ such that $U=P_{j}(G)\leq_{\mathrm{o}}G$ is uniformly powerful and forms the starting point for the analysis carried out there. In particular, there are matching filtration series
@@
Hausdorff dimensions of subgroups in $G$ with respect to the lower $p$-series $\mathcal{L}$ are closely linked to Hausdorff dimensions of subgroups in $U$ with respect to $\mathcal{S}_{U}$. Indeed, for every $H\leq_{\mathrm{c}}G$ and every open subgroup $W\leq_{\mathrm{o}}H$ with $W\subseteq U$,
@@
Moreover, we claim that, if $W$ is chosen so that it is powerful (using [5, Prop. 3.9]) with corresponding Lie sublattice $M\leq_{\mathrm{o}}L$, then
@@
To see (3.3), let $i\in\mathbb{N}$ with $i\geq j$ and choose $m\in\mathbb{N}$ such that $U^{p^{m}}\leq U_{i^{*}}$, and accordingly $p^{m}L\leq L_{i^{*}}$. By [5, Cor. 4.15], multiplicative cosets of $U^{p^{m}}$ in $U$ correspond to additive cosets of $p^{m}L$ in $L$. Consequently, $U_{i^{*}}$ is the union of as many $U^{p^{m}}$-cosets as $L_{i^{*}}$ is the union of $p^{m}L$-cosets, hence
@@
$[x_{j,k},y]=\begin{cases}x_{j,k+1}&\text{if $1\leq k\leq q_{j}-1$,}\\ x_{j,1}^{\,p}&\text{if $k=q_{j}$.}\end{cases}$
@@
$(1,\,\nicefrac{{1}}{{2}},\nicefrac{{1}}{{2}},\,\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}},\nicefrac{{1}}{{3}},\,\ldots,\,\underbrace{\nicefrac{{1}}{{q_{m}}},\ldots,\nicefrac{{1}}{{q_{m}}}}_{\text{$q_{m}$ entries}}).$
@@
We arrive at a lower bound as follows. For $H\leq_{\mathrm{c}}G$ of the form $H=\prod_{j=1}^{m}B_{j}$ with $B_{j}\leq A_{j}$ and $\dim(B_{j})=d_{j}$, say, a direct analysis, similar to the one in the proof of Theorem 3.1, yields
@@
as $m$ tends to infinity; compare with [20, Cor. 1] (and [2] for a more careful approximation). These estimates describe the size of the Hausdorff spectrum asymptotically and imply (3.4). ∎

### v3

@@ 2. Stratifications of the lower $p$-series
$P_{i-2cd}(G),P_{i-2cd}(G)]\subseteq P_{2i-4cd}(G)\subseteq P_{i+2cd}(G).$
In conjunction with [6, Cor. 3.5], the description of the Lie bracket in terms of the group multiplication, as described in [5, Sec. 4] and recorded in [16, Equ. (2.1)] yields the corresponding result on the Lie lattice side,
$L_{i^{*}-2cd},L_{i^{*}-2cd}]_{\mathrm{Lie}}\subseteq L_{i^{*}+2cd},$
where we write $\cdot,\cdot]_{\mathrm{Lie}}$ for the Lie bracket to distinguish it from group commutators. Similar to the simpler situation discussed above, the finite abelian group section $P_{i-2cd}(G)/P_{i+2cd}(G)$ of $G$ corresponds to a finite abelian Lie ring section $L_{i^{*}-2cd}/L_{i^{*}+2cd}$ of the $\mathbb{Z}_{p}$-Lie lattice $L$: the multiplicative cosets are the same as the additive cosets, multiplication and addition are the same and the Lie bracket is trivial; compare [8, Thm. 4.5].
Indeed, we already observed that the Lie bracket on $L_{i^{*}-2cd}$ is trivial modulo $L_{i^{*}+2cd}$, and the Lie ring $L_{i^{*}-2cd}/L_{i^{*}+2cd}$ is abelian. In conjunction with [16, Lem. B.2], the description of group multiplication via the Hausdorff series $\Phi$, namely
@@
where $\mu_{U}$ denotes the normalised Haar measure on the group $U$. We recall that the normalised Haar measures $\mu_{U}$ and $\mu_{L}$ on the compact space $U=L$ are the same; see [16, Prop. A.2] and [4, Lem. 1.9.4]. This gives
@@
The uniqueness of the parameters $\xi_{1},\ldots,\xi_{d}$ follows by looking at the elementary divisors of increasing sections, similar to the proof of Lemma 2.1. Indeed, suppose that $j$, $U$, $\xi_{1},\ldots,\xi_{d}$, $c$ and $\check{U}_{i}$ satisfy the assertions of Theorem 1.2. Let $L$ be the $\mathbb{Z}_{p}$-Lie lattice associated to $U$. By [5, Prop. 3.9], there is $j^{\prime}\in\mathbb{N}$ with $j^{\prime}\geq j$ such that, for all $i\in\mathbb{N}$ with $i\geq j^{\prime}$, the subgroups $P_{i}(G),\check{U}_{i}\trianglelefteq_{\mathrm{o}}G$ are uniformly powerful and thus give rise to full $\mathbb{Z}_{p}$-sublattices $A_{i},\check{M}_{i}\subseteq_{\mathrm{o}}\mathbb{Q}_{p}\otimes_{\mathbb{Z}_{p}}L$, by the Lie correspondence.
@@ 3. Hausdorff spectra with respect to the lower $p$-series
Furthermore, every subgroup $H\subseteq_{\mathrm{c}}L$ has strong Hausdorff dimension with respect to $\mathcal{S}$.
@@
We write $\mathcal{S}\colon L=L_{0}\supseteq L_{1}\supseteq\ldots$ and choose a frame $(x_{1},\ldots,x_{d})$ for a stratification of $\mathcal{S}$. We also recall that $0<\xi_{1}\leq\xi_{2}\leq\ldots\leq\xi_{d}$. By [18, Lem. 2.2], any two equivalent filtration series define the same Hausdorff dimension on closed subgroups of $L$. Hence we may suppose that
@@
Let $H\subseteq_{\mathrm{c}}L$ be a $\mathbb{Z}_{p}$-sublattice of dimension $\dim_{\mathbb{Z}_{p}}(H)=r$, say. Since $\mathbb{Z}_{p}$ is a principal ideal domain, $H$ has a $\mathbb{Z}_{p}$-basis $(y_{1},\ldots,y_{r})$ such that the parameters
@@
It suffices to prove that the upper limit on the right-hand side equals 0; in particular, this also shows that it is a proper limit. Since $H_{2}=\mathbb{Z}_{p}y_{r}$, we find that for each $i\in\mathbb{N}$ there exists $n_{i}\in\mathbb{N}_{0}$ such that
@@
As in the proof of Theorem 1.2 we find $j\in\mathbb{N}$ such that the subgroup $U=P_{j}(G)\subseteq_{\mathrm{o}}G$ is uniformly powerful and forms the starting point for the analysis carried out there. In particular, there are matching filtration series
@@
Hausdorff dimensions of subgroups in $G$ with respect to the lower $p$-series $\mathcal{L}$ are closely linked to Hausdorff dimensions of subgroups in $U$ with respect to $\mathcal{S}_{U}$. Indeed, for every subgroup $H\subseteq_{\mathrm{c}}G$ and every open subgroup $W\subseteq_{\mathrm{o}}H$ with $W\subseteq U$,
@@
Moreover, we claim that, if $W$ is chosen so that it is powerful (using [5, Prop. 3.9]) with corresponding Lie sublattice $M\subseteq_{\mathrm{o}}L$, then
@@
To see (3.3), let $i\in\mathbb{N}$ with $i\geq j$ and choose $m\in\mathbb{N}$ such that $U^{p^{m}}\subseteq U_{i^{*}}$, and accordingly $p^{m}L\subseteq L_{i^{*}}$. By [5, Cor. 4.15], multiplicative cosets of $U^{p^{m}}$ in $U$ correspond to additive cosets of $p^{m}L$ in $L$. Consequently, $U_{i^{*}}$ is the union of as many $U^{p^{m}}$-cosets as $L_{i^{*}}$ is the union of $p^{m}L$-cosets, hence
@@
$x_{j,k},y]=\begin{cases}x_{j,k+1}&\text{if $1\leq k\leq q_{j}-1$,}\\ x_{j,1}^{\,p}&\text{if $k=q_{j}$.}\end{cases}$
@@
[…]
@@
We arrive at a lower bound as follows. For every subgroup $H\subseteq_{\mathrm{c}}G$ of the form $H=\prod_{j=1}^{m}B_{j}$ with $B_{j}\subseteq A_{j}$ and $\dim(B_{j})=d_{j}$, say, a direct analysis, similar to the one in the proof of Theorem 3.1, yields
@@
as $m$ tends to infinity; compare with [21, Cor. 1] (and [2] for a more careful approximation). These estimates describe the size of the Hausdorff spectrum asymptotically and imply (3.4). ∎